OpenStudy (anonymous):

An analysis of a sugar compound is found to contain 132 grams of carbon (C), 22 grams of hydrogen (H), and 176 grams of oxygen (O). Determine the empirical formula of the compound. CHO C11H22O11 C6H12O6 CH2O

OpenStudy (anonymous):

@Astrophysics

OpenStudy (♪chibiterasu):

You have t convert to moles first. After that, divide each value by the smallest value that was calculated when converting. Round and use that to determine the empirical formula.

OpenStudy (anonymous):

can you help me solve it

OpenStudy (aaronq):

The only formula you need is: \(\sf moles=\dfrac{mass}{Molar~mass}\) for example, for hydrogen, \(\sf moles_H=\dfrac{mass_H}{Molar~mass_H}=\dfrac{22~g}{1.008~g/mol}\approx 22~moles \)